# ATLAS: Adaptive Topology-based Learning at Scale for Homophilic and Heterophilic Graphs

### v1

@@ 1 Introduction
Node classification, a fundamental problem in graph learning, involves identifying labels of nodes in a graph and has wide applications in many domains including social networks, citation networks, recommendation systems, knowledge graphs and bioinformatics  (Khemani2024GNNReview; Wu2019Survey; Zhou2021GNNReview). Accurate classification requires two complementary pieces of information–(i) the features at each node, and (ii) the connections between the node and its neighbors. Neural network methods such as Multi-Layer Perceptrons (MLPs) are fast but do not include information about the connections. Graph Neural Networks (GNNs) address this problem by aggregating the features between neighboring nodes, but the process is expensive, and difficult to scale to large graphs. Although the graph structure can be represented as feature vectors using different node embedding techniques (Perozzi2014DeepWalk; Grover2016node2vec; Tang2015LINE), or through the use of community detection (Sun2019vGraph; Kaminski2024CommunityFeatures), the issue remains as to how many hops of neighbors should be considered and how fine-grained the communities should be. Larger hops or coarse grained community can lead to information smoothing, while smaller hops or fine grained communities can lead to information loss. Further, the hypothesis that aggregating features of neighbors can improve accuracy of node classification is only true for homophilic networks (where nodes of similar classes are connected). In heterophilic networks, where the connection between nodes need not imply similarity of class, this strategy leads to lower accuracy. Based on these observations, we posit, matching structural information (i.e. size of hops or communities) with how well it aligns with the classification is necessary for producing accurate results.
@@ 1.1 Related Work
Graph Neural Networks (GNNs) have become a core tool for learning on graphs (kipf2017semi; hamilton2017inductive). Most algorithms follow a message-passing paradigm, aggregating transformed neighbor features into topology-aware embeddings, which implicitly assumes homophily (wu2019simplifying). The same bias can blur informative distinctions on weakly homophilous or heterophilous graphs (zhu2020beyond; platonov2023characterizing).
@@
Scaling GNNs on large graphs. Scaling GNNs on large graphs is challenging due to memory and aggregation costs. Sampling-based methods approximate full-batch propagation using node-, layer-, or subgraph-level sampling (GraphSAGE, FastGCN, Cluster-GCN, GraphSAINT, LABOR) (hamilton2017inductive; chen2018fastgcn; chiang2019cluster; Zeng2020GraphSAINT; balin2023labor), but introduce stochasticity that affects convergence and reproducibility (chen2018fastgcn; zou2019layer). Decoupled models instead precompute feature diffusion and train MLPs on fixed graph-derived features, enabling i.i.d. node mini-batching and fast inference (SGC, SIGN, SAGN, GAMLP, SCARA, LD2) (wu2019simplifying; rossi2020sign; sun2021sagn; chien2022gamlp; liao2022scara; liao2023ld2).
@@
Learning on non-homophilous graphs. For non-homophilous graphs, one line of work preserves self-features while carefully injecting neighborhood information (H2GCN, GloGNN) (zhu2020beyond; li2022glognn), or reweights neighbors to downweight harmful edges (GPR-GNN, FAGCN) (chien2022adaptive; bo2021fagcn). Others exploit higher-order propagation or spectral filters to capture both homophilic and heterophilic signals (MixHop, JacobiConv, BernNet, GBK-GNN) (abu-el-haija2019mixhop; wang2022jacobiconv; he2021bernnet; du2022gbkgnn). See zheng2022heterophily_survey; luan2024a_heterophilic_handbook for broader surveys.
@@
Community-aware node embeddings. Several works use community structure as an explicit representation for downstream prediction. Sun2019vGraph propose vGraph, a generative model that jointly infers discrete communities and continuous node embeddings by reconstructing edges, so that community assignments act as latent variables guiding representation learning. Closer to our setting, Kaminski2024CommunityFeatures construct community-aware node features (e.g., counts and statistics over community memberships in a node’s ego-network) and feed them into standard classifiers, showing that purely community-derived signals can already yield strong performance on node-level tasks.
@@
Graph–task alignment and community structure. A related line of work asks when a graph’s communities are informative for the labels and how this alignment controls the benefit of message passing. hussain2021interplay vary homophily and community structure in real graphs and define a measure of label–community correlation, showing that GNN gains are largest when labels follow communities and can vanish when they do not. This links the classical “cluster assumption” in semi-supervised learning (chapelle2006semisupervised) with recent analyses of graph–task and NTK–graph alignment in GNN training dynamics (yang2024training), and motivates methods that treat communities as task-relevant structural signals.
@@
Community-guided graph rewiring. Building on modularity-based detection (Newman2006; Blondel2008), ComMa and ComFy (RubioMadrigal2025ComFy) use community structure and feature similarity to rewire intra- and inter-community edges, improving label–community alignment and GNN accuracy on both homophilic and heterophilic graphs.
@@ 1.2 Our Contribution
    Bridging Frameworks. Unlike prior MLP-based models designed primarily for heterophilic graphs, ATLAS effectively supports both homophilic and heterophilic settings, thereby minimizing the accuracy gap traditionally observed between MLPs and GNNs. Moreover, its high inference efficiency positions ATLAS as a practical and scalable alternative to GNNs.
@@ 3 Methodology
Feature Augmentation. For a given resolution parameter $\gamma$, let the communities be $\mathbf{c}^{(\gamma)}\in\{1,\dots,k_{\gamma}\}$, and and each node is assigned to one of the communities in $k_{\gamma}$. This assignment is represented as a one-hot encoded matrix $\mathbf{H}^{(\gamma)}$( equation 3). To reduce dimensionality, each one-hot matrix is projected into a dense embedding space using a trainable weight matrix $\mathbf{W}^{(\gamma)}$ ( equation 4). The embeddings from all resolutions are concatenated to form $\mathbf{E}$ ( equation 5), which is then further concatenated with the original features $\mathbf{X}$ to yield the augmented feature matrix $\mathbf{Z}$. The augmented feature matrix $\mathbf{Z}$ is fed to an MLP $f_{\theta}$ to produce logits; a task-dependent function $\phi$ (e.g., softmax or elementwise sigmoid), applied row-wise, converts them to probabilities $\hat{\mathbf{Y}}$ ( equation 7).
@@ 4 Empirical Evaluation
Q1 (homophily–heterophily regime). Homophilic: GCN (kipf2017semi), GraphSAGE (hamilton2017inductive), GAT (velickovic2018graph). Heterophily-oriented: H2GCN (zhu2020beyond), LinkX (lim2021linkx), GPR-GNN (chien2022adaptive), FSGNN (maurya2022fsgnn), GloGNN (li2022glognn), FAGCN (bo2021fagcn), GBK-GNN (du2022gbkgnn), JacobiConv (wang2022jacobiconv), ACM-GCN (luan2022revisiting), BernNet (he2021bernnet).
Q2 (scalability). Decoupled: SGC (wu2019simplifying), SIGN (rossi2020sign), SAGN (sun2021sagn), GAMLP (chien2022gamlp). Sampling-based: GraphSAGE (hamilton2017inductive), ClusterGCN (chiang2019cluster), GraphSAINT (Zeng2020GraphSAINT). Descriptions of these methods are provided in the Appendix.
@@ 4.1 Q1: Accuracy Across Homophily Regimes
Table 1: Eight-benchmark comparison across homophily regimes. Baseline heterophily-oriented model results are from platonov2023critical; luan2024b_reevaluating_heterophily. Bottom rows report ATLAS improvements over baselines (absolute percentage points). Cells highlighted in yellow indicate the best score for each dataset.
@@
MLP(2L) 75.44 ± 1.97 72.97 ± 0.90 87.25 ± 0.41 36.00 ± 4.69 39.83 ± 0.48 34.96 ± 0.71 34.29 ± 3.34 65.58 ± 0.34
GCN 87.01 ± 1.04 74.93 ± 1.32 86.71 ± 0.42 37.11 ± 3.04 42.78 ± 0.14 28.49 ± 0.91 32.70 ± 1.73 45.68 ± 0.38
SAGE 87.50 ± 0.87 80.95 ± 0.92 88.42 ± 0.55 38.83 ± 4.26 44.67 ± 0.51 34.08 ± 1.07 33.32 ± 1.75 76.21 ± 0.65
GAT 87.74 ± 0.88 75.31 ± 1.35 86.18 ± 0.64 37.18 ± 3.44 43.25 ± 0.85 29.11 ± 1.23 32.61 ± 2.06 47.16 ± 0.66
H2GCN 87.52 ± 0.61 73.35 ± 1.01 87.78 ± 0.28 26.75 ± 3.64 36.47 ± 0.23 38.85 ± 1.17 35.10 ± 1.15 60.11 ± 0.52
LinkX 82.62 ± 1.44 81.15 ± 1.23 88.12 ± 0.47 40.10 ± 2.21 52.66 ± 0.64 35.64 ± 1.36 42.34 ± 4.13 56.15 ± 0.93
GPR-GNN 79.51 ± 0.36 72.94 ± 0.97 85.07 ± 0.09 39.93 ± 3.30 44.88 ± 0.34 39.30 ± 0.27 38.95 ± 1.99 64.85 ± 0.27
FSGNN 87.51 ± 1.21 82.76 ± 0.61 90.11 ± 0.43 40.61 ± 2.97 52.74 ± 0.83 37.65 ± 0.79 35.92 ± 1.32 79.92 ± 0.56
GloGNN 87.67 ± 1.16 73.39 ± 1.17 90.32 ± 0.54 25.90 ± 3.58 36.89 ± 0.14 39.65 ± 1.03 35.11 ± 1.24 59.63 ± 0.69
FAGCN 88.85 ± 1.36 77.75 ± 1.05 89.98 ± 0.54 41.90 ± 2.72 44.12 ± 0.30 31.59 ± 1.37 41.08 ± 2.27 65.22 ± 0.56
GBK-GNN 87.09 ± 1.52 81.01 ± 0.67 88.88 ± 0.44 39.61 ± 2.60 45.98 ± 0.71 38.47 ± 1.53 35.51 ± 1.65 74.57 ± 0.47
JacobiConv 89.61 ± 0.96 68.66 ± 0.65 89.99 ± 0.39 39.00 ± 4.20 43.55 ± 0.48 37.48 ± 0.76 29.71 ± 1.66 71.14 ± 0.42
BernNet 88.52 ± 0.95 77.00 ± 0.65 88.48 ± 0.41 40.90 ± 4.06 44.64 ± 0.56 41.79 ± 1.01 41.18 ± 1.77 65.56 ± 1.34
ACM-GCN 89.75 ± 1.16 74.95 ± 1.16 90.96 ± 0.62 42.73 ± 3.59 52.49 ± 0.24 41.86 ± 1.48 42.35 ± 1.97 71.89 ± 0.61
ATLAS 87.09 ± 1.62 82.19 ± 0.73 88.85 ± 0.48 42.76 ± 3.47 53.15 ± 0.61 38.48 ± 0.93 40.35 ± 1.53 66.22 ± 0.53
ATLAS-NF 86.73 ± 1.04 83.02 ± 0.74 88.76 ± 0.36 40.02 ± 2.79 52.30 ± 0.64 34.26 ± 0.97 36.98 ± 2.37 77.94 ± 0.48
ATLAS–MLP (pp) +11.65 +9.22 +1.60 +6.76 +13.32 +3.52 +6.06 +0.64
ATLAS–GCN (pp) +0.08 +7.26 +2.14 +5.65 +10.37 +9.99 +7.65 +20.54
ATLAS–Average (pp) +0.92 +5.97 +0.40 +5.15 +8.51 +2.13 +3.91 +1.67
Table 2: Large-graph performance. Baselines from Zeng2020GraphSAINT; hu2020ogb. Bottom rows report ATLAS improvements over MLP and over GCN (absolute units). Cells highlighted in yellow indicate the best score for each dataset.
@@
MLP 0.7435 ± 0.0016 0.6106 ± 0.0008 0.4717 ± 0.0011 0.6546 ± 0.0011 0.8204 ± 0.0002
GCN 0.9330 ± 0.0001 0.7564 ± 0.0021 0.4920 ± 0.0030 0.3780 ± 0.0010 0.2810 ± 0.0050
GraphSAGE 0.9530 ± 0.0010 0.8061 ± 0.0016 0.5010 ± 0.0130 0.6340 ± 0.0060 0.7580 ± 0.0020
ClusterGCN 0.9540 ± 0.0010 0.7862 ± 0.0061 0.4810 ± 0.0050 0.6090 ± 0.0050 0.7590 ± 0.0080
GraphSAINT 0.9660 ± 0.0010 0.7536 ± 0.0034 0.5110 ± 0.0010 0.6530 ± 0.0030 0.8150 ± 0.0010
SGC 0.9351 ± 0.0004 0.6748 ± 0.0011 0.5035 ± 0.0005 0.2356 ± 0.0002 0.2262 ± 0.0028
SIGN 0.9595 ± 0.0002 0.8052 ± 0.0016 0.5160 ± 0.0011 0.5798 ± 0.0012 0.7424 ± 0.0002
SAGN 0.9648 ± 0.0003 0.8121 ± 0.0007 0.5007 ± 0.0011 0.6155 ± 0.0040 0.7682 ± 0.0115
GAMLP 0.9673 ± 0.0003 0.8376 ± 0.0019 0.5258 ± 0.0012 0.5784 ± 0.0154 0.7599 ± 0.0026
ATLAS 0.9574 ± 0.0004 0.7865 ± 0.0053 0.5104 ± 0.0039 0.6546 ± 0.0011 0.8204 ± 0.0002
ATLAS-NF 0.9410 ± 0.0007 0.7507 ± 0.0030 0.5201 ± 0.0012 0.5740 ± 0.0021 0.7783 ± 0.0010
ATLAS–MLP +0.2139 +0.1759 +0.0387 +0.0000 +0.0000
ATLAS–GCN +0.0244 +0.0301 +0.0184 +0.2766 +0.5394
ATLAS–Average +0.0267 +0.0262 +0.0101 +0.1059 +0.1615

### v2

@@ 1 Introduction
Node classification, a fundamental problem in graph learning, involves identifying labels of nodes in a graph and has wide applications in many domains including social networks, citation networks, recommendation systems, knowledge graphs and bioinformatics  (Khemani, 2024; Wu et al., 2019b; Zhou et al., 2021). Accurate classification requires two complementary pieces of information–(i) the features at each node, and (ii) the connections between the node and its neighbors. Neural network methods such as Multi-Layer Perceptrons (MLPs) are fast but do not include information about the connections. Graph Neural Networks (GNNs) address this problem by aggregating the features between neighboring nodes, but the process is expensive, and difficult to scale to large graphs. Although the graph structure can be represented as feature vectors using different node embedding techniques (Perozzi et al., 2014; Grover and Leskovec, 2016; Tang et al., 2015), or through the use of community detection (Sun et al., 2019; Kamiński et al., 2024), the issue remains as to how many hops of neighbors should be considered and how fine-grained the communities should be. Larger hops or coarse grained community can lead to information smoothing, while smaller hops or fine grained communities can lead to information loss. Further, the hypothesis that aggregating features of neighbors can improve accuracy of node classification is only true for homophilic networks (where nodes of similar classes are connected). In heterophilic networks, where the connection between nodes need not imply similarity of class, this strategy leads to lower accuracy. Based on these observations, we posit, matching structural information (i.e. size of hops or communities) with how well it aligns with the classification is necessary for producing accurate results.
@@ 1.1 Related Work
Graph Neural Networks (GNNs) have become a core tool for learning on graphs (Kipf and Welling, 2017; Hamilton et al., 2017). Most algorithms follow a message-passing paradigm, aggregating transformed neighbor features into topology-aware embeddings, which implicitly assumes homophily (Wu et al., 2019a). The same bias can blur informative distinctions on weakly homophilous or heterophilous graphs (Zhu et al., 2020; Platonov et al., 2023a).
@@
Scaling GNNs on large graphs. Scaling GNNs on large graphs is challenging due to memory and aggregation costs. Sampling-based methods approximate full-batch propagation using node-, layer-, or subgraph-level sampling (GraphSAGE, FastGCN, Cluster-GCN, GraphSAINT, LABOR) (Hamilton et al., 2017; Chen et al., 2018b; Chiang et al., 2019; Zeng et al., 2020; Balın and Çatalyürek, 2023), but introduce stochasticity that affects convergence and reproducibility (Chen et al., 2018b; Zou et al., 2019). Decoupled models instead precompute feature diffusion and train MLPs on fixed graph-derived features, enabling i.i.d. node mini-batching and fast inference (SGC, SIGN, SAGN, GAMLP, SCARA, LD2) (Wu et al., 2019a; Rossi et al., 2020; Sun et al., 2021; Chien et al., 2022; Liao et al., 2022; 2023).
@@
Learning on non-homophilous graphs. For non-homophilous graphs, one line of work preserves self-features while carefully injecting neighborhood information (H2GCN, GloGNN) (Zhu et al., 2020; Li et al., 2022), or reweights neighbors to downweight harmful edges (GPR-GNN, FAGCN) (Chien et al., 2021; Bo et al., 2021). Others exploit higher-order propagation or spectral filters to capture both homophilic and heterophilic signals (MixHop, JacobiConv, BernNet, GBK-GNN) (Abu-El-Haija et al., 2019; Wang and Zhang, 2022; He et al., 2021; Du et al., 2022). See Zheng et al. (2022); Luan et al. (2024b) for broader surveys.
@@
Community-aware node embeddings. Several works use community structure as an explicit representation for downstream prediction. Sun et al. (2019) propose vGraph, a generative model that jointly infers discrete communities and continuous node embeddings by reconstructing edges, so that community assignments act as latent variables guiding representation learning. Closer to our setting, Kamiński et al. (2024) construct community-aware node features (e.g., counts and statistics over community memberships in a node’s ego-network) and feed them into standard classifiers, showing that purely community-derived signals can already yield strong performance on node-level tasks.
@@
Graph–task alignment and community structure. A related line of work asks when a graph’s communities are informative for the labels and how this alignment controls the benefit of message passing. Hussain et al. (2021) vary homophily and community structure in real graphs and define a measure of label–community correlation, showing that GNN gains are largest when labels follow communities and can vanish when they do not. This links the classical “cluster assumption” in semi-supervised learning (Chapelle et al., 2006) with recent analyses of graph–task and NTK–graph alignment in GNN training dynamics (Yang et al., 2024), and motivates methods that treat communities as task-relevant structural signals.
@@
Community-guided graph rewiring. Building on modularity-based detection (Newman, 2006; Blondel et al., 2008), ComMa and ComFy (Rubio-Madrigal et al., 2025) use community structure and feature similarity to rewire intra- and inter-community edges, improving label–community alignment and GNN accuracy on both homophilic and heterophilic graphs.
@@ 1.2 Our Contribution
    Analysis. We study how performance varies with community refinement and show consistent trends across “high/low/negative structural-bias” regimes, offering an interpretable view of when topology helps (Section  5).
  5. 5.
    Bridging Frameworks. Unlike prior MLP-based models designed primarily for heterophilic graphs, ATLAS effectively supports both homophilic and heterophilic settings, thereby minimizing the accuracy gap traditionally observed between MLPs and GNNs.
@@ 3 Methodology
Feature Augmentation. For a given resolution parameter $\gamma$, let the communities be $\mathbf{c}^{(\gamma)}\in\{1,\dots,k_{\gamma}\}$, and each node is assigned to one of the communities in $k_{\gamma}$. This assignment is represented as a one-hot encoded matrix $\mathbf{H}^{(\gamma)}$( equation 3). To reduce dimensionality, each one-hot matrix is projected into a dense embedding space using a trainable weight matrix $\mathbf{W}^{(\gamma)}$ ( equation 4). The embeddings from all resolutions are concatenated to form $\mathbf{E}$ ( equation 5), which is then further concatenated with the original features $\mathbf{X}$ to yield the augmented feature matrix $\mathbf{Z}$. The augmented feature matrix $\mathbf{Z}$ is fed to an MLP $f_{\theta}$ to produce logits; a task-dependent function $\phi$ (e.g., softmax or elementwise sigmoid), applied row-wise, converts them to probabilities $\hat{\mathbf{Y}}$ ( equation 7).
@@ 4 Empirical Evaluation
Q1 (homophily–heterophily regime). Homophilic: GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018). Heterophily-oriented: H2GCN (Zhu et al., 2020), LinkX (Lim et al., 2021), GPR-GNN (Chien et al., 2021), FSGNN (Maurya et al., 2022), GloGNN (Li et al., 2022), FAGCN (Bo et al., 2021), GBK-GNN (Du et al., 2022), JacobiConv (Wang and Zhang, 2022), ACM-GCN (Luan et al., 2022), BernNet (He et al., 2021).
Q2 (scalability). Decoupled: SGC (Wu et al., 2019a), SIGN (Rossi et al., 2020), SAGN (Sun et al., 2021), GAMLP (Chien et al., 2022). Sampling-based: GraphSAGE (Hamilton et al., 2017), ClusterGCN (Chiang et al., 2019), GraphSAINT (Zeng et al., 2020). Descriptions of these methods are provided in the Appendix.
@@ 4.1 Q1: Accuracy Across Homophily Regimes
Table 1: Eight-benchmark comparison across homophily regimes. Baseline heterophily-oriented model results are from Platonov et al. (2023b); Luan et al. (2024a). Bottom rows report ATLAS improvements over baselines (absolute percentage points). Cells highlighted in yellow indicate the best score for each dataset.
@@
MLP(2L) 75.44 $\pm$ 1.97 72.97 $\pm$ 0.90 87.25 $\pm$ 0.41 36.00 $\pm$ 4.69 39.83 $\pm$ 0.48 34.96 $\pm$ 0.71 34.29 $\pm$ 3.34 65.58 $\pm$ 0.34
GCN 87.01 $\pm$ 1.04 74.93 $\pm$ 1.32 86.71 $\pm$ 0.42 37.11 $\pm$ 3.04 42.78 $\pm$ 0.14 28.49 $\pm$ 0.91 32.70 $\pm$ 1.73 45.68 $\pm$ 0.38
SAGE 87.50 $\pm$ 0.87 80.95 $\pm$ 0.92 88.42 $\pm$ 0.55 38.83 $\pm$ 4.26 44.67 $\pm$ 0.51 34.08 $\pm$ 1.07 33.32 $\pm$ 1.75 76.21 $\pm$ 0.65
GAT 87.74 $\pm$ 0.88 75.31 $\pm$ 1.35 86.18 $\pm$ 0.64 37.18 $\pm$ 3.44 43.25 $\pm$ 0.85 29.11 $\pm$ 1.23 32.61 $\pm$ 2.06 47.16 $\pm$ 0.66
H2GCN 87.52 $\pm$ 0.61 73.35 $\pm$ 1.01 87.78 $\pm$ 0.28 26.75 $\pm$ 3.64 36.47 $\pm$ 0.23 38.85 $\pm$ 1.17 35.10 $\pm$ 1.15 60.11 $\pm$ 0.52
LinkX 82.62 $\pm$ 1.44 81.15 $\pm$ 1.23 88.12 $\pm$ 0.47 42.34 $\pm$ 4.13 52.66 $\pm$ 0.64 35.64 $\pm$ 1.36 40.10 $\pm$ 2.21 56.15 $\pm$ 0.93
GPR-GNN 79.51 $\pm$ 0.36 72.94 $\pm$ 0.97 85.07 $\pm$ 0.09 39.93 $\pm$ 3.30 44.88 $\pm$ 0.34 39.30 $\pm$ 0.27 38.95 $\pm$ 1.99 64.85 $\pm$ 0.27
FSGNN 87.51 $\pm$ 1.21 82.76 $\pm$ 0.61 90.11 $\pm$ 0.43 40.61 $\pm$ 2.97 52.74 $\pm$ 0.83 37.65 $\pm$ 0.79 35.92 $\pm$ 1.32 79.92 $\pm$ 0.56
GloGNN 87.67 $\pm$ 1.16 73.39 $\pm$ 1.17 90.32 $\pm$ 0.54 25.90 $\pm$ 3.58 36.89 $\pm$ 0.14 39.65 $\pm$ 1.03 35.11 $\pm$ 1.24 59.63 $\pm$ 0.69
FAGCN 88.85 $\pm$ 1.36 77.75 $\pm$ 1.05 89.98 $\pm$ 0.54 41.90 $\pm$ 2.72 44.12 $\pm$ 0.30 31.59 $\pm$ 1.37 41.08 $\pm$ 2.27 65.22 $\pm$ 0.56
GBK-GNN 87.09 $\pm$ 1.52 81.01 $\pm$ 0.67 88.88 $\pm$ 0.44 39.61 $\pm$ 2.60 45.98 $\pm$ 0.71 38.47 $\pm$ 1.53 35.51 $\pm$ 1.65 74.57 $\pm$ 0.47
JacobiConv 89.61 $\pm$ 0.96 68.66 $\pm$ 0.65 89.99 $\pm$ 0.39 39.00 $\pm$ 4.20 43.55 $\pm$ 0.48 37.48 $\pm$ 0.76 29.71 $\pm$ 1.66 71.14 $\pm$ 0.42
BernNet 88.52 $\pm$ 0.95 77.00 $\pm$ 0.65 88.48 $\pm$ 0.41 40.90 $\pm$ 4.06 44.64 $\pm$ 0.56 41.79 $\pm$ 1.01 41.18 $\pm$ 1.77 65.56 $\pm$ 1.34
ACM-GCN 89.75 $\pm$ 1.16 74.95 $\pm$ 1.16 90.96 $\pm$ 0.62 42.73 $\pm$ 3.59 52.49 $\pm$ 0.24 41.86 $\pm$ 1.48 42.35 $\pm$ 1.97 71.89 $\pm$ 0.61
ATLAS 87.09 $\pm$ 1.62 82.19 $\pm$ 0.73 88.85 $\pm$ 0.48 42.76 $\pm$ 3.47 53.15 $\pm$ 0.61 38.48 $\pm$ 0.93 40.82 $\pm$ 2.12 66.22 $\pm$ 0.53
ATLAS-NF 86.73 $\pm$ 1.04 83.02 $\pm$ 0.74 88.76 $\pm$ 0.36 40.02 $\pm$ 2.79 52.30 $\pm$ 0.64 34.26 $\pm$ 0.97 36.98 $\pm$ 2.37 78.10 $\pm$ 0.59
ATLAS–MLP (pp) +11.65 +9.22 +1.60 +6.76 +13.32 +3.52 +6.53 +0.64
ATLAS–GCN (pp) +0.08 +7.26 +2.14 +5.65 +10.37 +9.99 +8.12 +20.54
ATLAS–Average (pp) +0.92 +5.97 +0.40 +4.99 +8.51 +2.13 +4.54 +1.67
Table 2: Large-graph performance. Baselines from Zeng et al. (2020); Hu et al. (2020). Bottom rows report ATLAS improvements over MLP and over GCN (absolute units). Cells highlighted in yellow indicate the best score for each dataset.
@@
MLP 0.7435 $\pm$ 0.0016 0.6106 $\pm$ 0.0008 0.4717 $\pm$ 0.0011 0.6546 $\pm$ 0.0011 0.8204 $\pm$ 0.0002
GCN 0.9330 $\pm$ 0.0001 0.7564 $\pm$ 0.0021 0.4920 $\pm$ 0.0030 0.3780 $\pm$ 0.0010 0.2810 $\pm$ 0.0050
GraphSAGE 0.9530 $\pm$ 0.0010 0.8061 $\pm$ 0.0016 0.5010 $\pm$ 0.0130 0.6340 $\pm$ 0.0060 0.7580 $\pm$ 0.0020
ClusterGCN 0.9540 $\pm$ 0.0010 0.7862 $\pm$ 0.0061 0.4810 $\pm$ 0.0050 0.6090 $\pm$ 0.0050 0.7590 $\pm$ 0.0080
GraphSAINT 0.9660 $\pm$ 0.0010 0.7536 $\pm$ 0.0034 0.5110 $\pm$ 0.0010 0.6530 $\pm$ 0.0030 0.8150 $\pm$ 0.0010
SGC 0.9351 $\pm$ 0.0004 0.6748 $\pm$ 0.0011 0.5035 $\pm$ 0.0005 0.2356 $\pm$ 0.0002 0.2262 $\pm$ 0.0028
SIGN 0.9595 $\pm$ 0.0002 0.8052 $\pm$ 0.0016 0.5160 $\pm$ 0.0011 0.5798 $\pm$ 0.0012 0.7424 $\pm$ 0.0002
SAGN 0.9648 $\pm$ 0.0003 0.8121 $\pm$ 0.0007 0.5007 $\pm$ 0.0011 0.6155 $\pm$ 0.0040 0.7682 $\pm$ 0.0115
GAMLP 0.9673 $\pm$ 0.0003 0.8376 $\pm$ 0.0019 0.5258 $\pm$ 0.0012 0.5784 $\pm$ 0.0154 0.7599 $\pm$ 0.0026
ATLAS 0.9574 $\pm$ 0.0004 0.7998 $\pm$ 0.0014 0.5156 $\pm$ 0.0014 0.6546 $\pm$ 0.0011 0.8204 $\pm$ 0.0002
ATLAS-NF 0.9410 $\pm$ 0.0007 0.7507 $\pm$ 0.0030 0.5272 $\pm$ 0.0011 0.5740 $\pm$ 0.0021 0.7783 $\pm$ 0.0010
ATLAS–MLP +0.2139 +0.1892 +0.0439 +0.0000 +0.0000
ATLAS–GCN +0.0244 +0.0434 +0.0236 +0.2766 +0.5394
ATLAS–Average +0.0267 +0.0395 +0.0153 +0.1059 +0.1615